# A Family of Berndt-Type Integrals and Associated Barnes Multiple Zeta Functions

### v1

@@ 5 Evaluations of Hyperbolic Summations via Jacobi Functions
$X_{p,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)=-(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p+1}\displaystyle\frac{\mathrm{P}_{p}\left(1-x\right)}{2}\sqrt{x(1-x)}$ (5.76)
@@
$\text{DB}_{p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.77)
@@
By Mathematica, we have
@@
Set $\Gamma=\Gamma(1/4)$ and $p=4m-1$. Then, for any integer $m>0$, we know that
@@
$\quad-\displaystyle\frac{1}{2^{4m+3}}(4m-2)\left(\mathrm{A}_{4m-2}\left(\displaystyle\frac{1}{2}\right)^{\prime}+\mathrm{A}_{4m-2}\left(\displaystyle\frac{1}{2}\right)\right)\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m+2}}{\pi^{(12m+3)/2}}.$ (5.78)
@@
Set $\Gamma=\Gamma(1/4)$. By Mathematica, we have
@@ 6 Evaluations of Berndt-type Integrals
With Mathematica, we can compute the following using the formulas from Theorem 6.1.
@@ 7 Berndt-type Integrals via Barnes Multiple Zeta Functions
In this section, we evaluate Barnes multiple zeta function through the results of Berndt-type integrals. Building upon the work of Bradshaw and Vignat [7], who evaluated generalized Berndt-type integrals (1.1) using the (alternating) Barnes multiple zeta function, we present key results. For instance, they established ([7, Prop. 2]):
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{a}dx}{(\cos x-\cosh x)^{b}}=2^{b}\Gamma(a+1)\zeta_{2b}(a+1,b|(1+i,1-i)^{b})\quad(a\geq 2b,b\geq 1)$ (7.89)
and
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{a}dx}{(\cos x+\cosh x)^{b}}=2^{b}\Gamma(a+1){\bar{\zeta}}_{2b}(a+1,b|(1+i,1-i)^{b})\quad(a\geq 0,b\geq 1),$ (7.90)
[…]
$\zeta_{N}(s,\omega|{\bf a}_{N}):=\displaystyle\sum_{n_{1}\geq 0,\ldots,n_{N}\geq 0}\displaystyle\frac{1}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}\quad(\Re(\omega)>0,\Re(s)>N)$ (7.91)
@@
Bradshaw and Vignat [7] provided integral representations for these functions, recovering a known result [15, Eq. (3.2)].
@@
Let $\Re(s)>N,\Re(w)>0$, and $\Re(a_{j})>0$ for $j=1,\ldots,N$. Then
$\zeta_{N}(s,w|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-wu}\prod_{j=1}^{N}(1-e^{-a_{j}u})^{-1}du.$
The alternating version satisfies
$\widetilde{\zeta}_{N}(s,w|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-wu}\prod_{j=1}^{N}(1+e^{-a_{j}u})^{-1}du.$
@@
Following [7, Prop. 2], we obtain
@@
$=2^{M+N}\Gamma\left(s\right)\zeta_{M+N}\left(s,w+\displaystyle\sum_{i=1}^{M}{a_{i}+}\displaystyle\sum_{i=1}^{N}{b_{i}}|\boldsymbol{c}_{M+N},\boldsymbol{\sigma}_{M+N}\right),$
@@
Applying this to our integral
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\sinh\left[\left(1+i\right)x\right]\sinh\left[\left(1-i\right)x\right]\cosh\left(\displaystyle\frac{1+i}{2}x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)}}$
$=4\Gamma\left(4m\right)\zeta_{4}\left(4m,3|\mathbf{c}_{4},\boldsymbol{\sigma}_{4}\right),$
@@
Employing the computational framework of Mathematica in conjunction with Theorem 7.2, we perform explicit calculations for integer values $m=1,2,3,4$.

### v3

@@ 5 Evaluations of Hyperbolic Summations via Jacobi Functions
$\tilde{G}_{p,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)=-(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p+1}\displaystyle\frac{\mathrm{P}_{p}\left(1-x\right)}{2}\sqrt{x(1-x)}$ (5.77)
@@
$G_{p,-1,2}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}G_{p-1,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}G_{p-1,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.78)
@@
By evaluating the right-hand side of (5.74) using Mathematica, we present several specific examples as follows:
@@
Setting p=4m1p=4m-1 and $x=1/2$ in Theorem 5.9, and noting that $\mathrm{P}_{4m-1}(1/2)=0$ (see the proof of Theorem 3.7 in [19]), we derive the following corollary.
@@
Set $\Gamma=\Gamma(1/4)$ and $\mathrm{A}_{p}^{\prime}(x):=\mathrm{d}\mathrm{A}_{p}(x)/\mathrm{d}x$. For any integer $m>0$, we have
@@
$\quad-\displaystyle\frac{1}{2^{4m+3}}\left(\mathrm{A}^{\prime}_{4m-2}\left(\displaystyle\frac{1}{2}\right)+\mathrm{A}_{4m-2}\left(\displaystyle\frac{1}{2}\right)\right)\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m+2}}{\pi^{(12m+3)/2}}.$ (5.79)
@@
Set $\Gamma=\Gamma(1/4)$. By Mathematica, we substitute $y=\pi$ and $x=1/2$ into Example $5.10$ and obtain the following identities
@@ 6 Evaluations of Berndt-type Integrals
Remark 6.2.
Note that $K({1}/{\sqrt{2}})=\displaystyle\frac{\Gamma^{2}(1/4)}{4\sqrt{\pi}}$ (see [8]). Consequently, the right-hand side of (6.1) can also be expressed in an alternative form:
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}$ $\in\mathbb{Q}\left[4K({1}/{\sqrt{2}})\right]^{4m-2}+\displaystyle\frac{\mathbb{Q}}{\sqrt{2}}\left[4K({1}/{\sqrt{2}})\right]^{4m-1}$
$\quad+\mathbb{Q}\left[4K({1}/{\sqrt{2}})\right]^{4m}+\displaystyle\frac{\mathbb{Q}}{\sqrt{2}}\displaystyle\frac{\left[4K({1}/{\sqrt{2}})\right]^{4m+1}}{\pi}$
$\quad+\mathbb{Q}\displaystyle\frac{\left[4K({1}/{\sqrt{2}})\right]^{4m+2}}{\pi^{2}}.$ (6.90)
The Gamma function form is preferred for our focus on numerical computation, special value correlations, and multiple zeta functions, while the $K$-function form is more appropriate for elliptic integral theory, result generality, and field-specific conventions. These expressions are provided for readers’ reference.
Using (6.85)-(6.89) and assisted by Mathematica, we compute the values of the five constant coefficients appearing on the right-hand side of the formula in Theorem 6.1 for $m=1,2,3,4$. This yields the following examples.
@@ 7 Berndt-type Integrals via Barnes Multiple Zeta Functions
In this section, we evaluate certain Barnes multiple zeta values via Berndt-type integrals. Building upon the work of Bradshaw and Vignat [9], who expressed generalized Berndt-type integrals (1.1) in terms of the (alternating) Barnes multiple zeta function, we record the following key identities. For instance, they proved [9, Prop. 2] that
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{a}\mathrm{d}x}{(\cos x-\cosh x)^{b}}=2^{b}\Gamma(a+1)\zeta_{2b}(a+1,b|(1+i,1-i)^{b})\quad(a\geq 2b,\ b\geq 1),$ (7.91)
$\displaystyle\int_{0}^{\infty}\displaystyle\frac{x^{a}\mathrm{d}x}{(\cos x+\cosh x)^{b}}=2^{b}\Gamma(a+1){\bar{\zeta}}_{2b}(a+1,b|(1+i,1-i)^{b})\quad(a\geq 0,\ b\geq 1).$ (7.92)
Here ${\boldsymbol{\sl{s}}}^{n}$ denotes the $n$-fold repetition of the string $\boldsymbol{\sl{s}}$}. Combined with the explicit evaluations of Xu and Zhao [24], these formulas yield closed-form expressions for the corresponding Barnes multiple zeta values (see [9, Cor. 2]).
For $a_{1},\ldots,a_{N}>0$ (write ${\bf a}_{N}:=\{a_{1},\ldots,a_{N}\}$), the Barnes multiple zeta function and its alternating analogue are defined by [3, 4, 9, 13] as
$\zeta_{N}(s,\omega|{\bf a}_{N}):=\displaystyle\sum_{n_{1}\geq 0,\ldots,n_{N}\geq 0}\displaystyle\frac{1}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}\quad(\Re(\omega)>0,\ \Re(s)>N),$ (7.93)
@@
Bradshaw and Vignat [9] also derived a convenient Laplace–Mellin formula for these series, recovering a known identity [17, Eq. (3.2)].
@@
Let $\Re(s)>N$, $\Re(\omega)>0$, and $\Re(a_{j})>0$ for $j=1,\ldots,N$. Then
$\zeta_{N}(s,\omega|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-\omega u}\prod_{j=1}^{N}(1-e^{-a_{j}u})^{-1}\,\mathrm{d}u,$
$\bar{\zeta}_{N}(s,\omega|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-\omega u}\prod_{j=1}^{N}(1+e^{-a_{j}u})^{-1}\,\mathrm{d}u.$
In a recent paper [27], the third author of this paper introduced the following generalized Barnes multiple zeta function:
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N}):=\displaystyle\sum_{n_{1}\geq 0,\ldots,n_{N}\geq 0}\displaystyle\frac{\sigma_{1}^{n_{1}}\cdots\sigma_{N}^{n_{N}}}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}\quad(\Re(\omega)>0,\ \Re(s)>N),$ (7.95)
where $\boldsymbol{\sigma}_{N}=(\sigma_{1},\ldots,\sigma_{N})\in\{\pm 1\}^{N}$. Moreover, Zhou [27, Prop. 5.9] also established the following integral representation:
Proposition 7.2.
Let $\Re(s)>N$, $\Re(\omega)>0$, and $\Re(a_{j})>0$ for $j=1,\ldots,N$. Then
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-\omega u}\prod_{j=1}^{N}\bigl(1-\sigma_{j}e^{-a_{j}u}\bigr)^{-1}\,\mathrm{d}u.$
@@
Following [9, Prop. 2], we obtain
@@
$=2^{M+N}\Gamma\left(s\right)\zeta_{M+N}\left(s,w+\displaystyle\sum_{i=1}^{M}{a_{i}+}\displaystyle\sum_{i=1}^{N}{b_{i}}|\boldsymbol{c}_{M+N};\boldsymbol{\sigma}_{M+N}\right),$
@@
By applying the product-to-sum formulas for trigonometric (and hyperbolic) functions, direct calculations yield
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\left[\cosh(2x)-\cos(2x)\right]\left[\cosh x+\cos x\right]}}=4\Gamma\left(4m\right)\zeta_{4}\left(4m,3|\mathbf{c}_{4},\boldsymbol{\sigma}_{4}\right)$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-1}\mathrm{d}x}{\sinh\left[\left(1+i\right)x\right]\sinh\left[\left(1-i\right)x\right]\cosh\left(\displaystyle\frac{1+i}{2}x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)}},$
@@
Employing the computational framework of Mathematica in conjunction with Theorem 7.3, we perform explicit calculations for $m=1,2,3,4$.
@@
Acknowledgments. The authors would like to thank the anonymous referees for their valuable comments and suggestions. Ce Xu is supported by the General Program of Natural Science Foundation of Anhui Province (Grant No. 2508085MA014).